A Company that makes shopping carts for supermarkets and other stores recently purchased some new equipment that reduces the labor content of the jobs needed to produce the shopping carts. Prior to buying the new equipment, the company used 4 workers, who produced an average of 82 carts per hour. Workers receive $11 per hour, and machine coast was $52 per hour. With the new equipment, it was possible to transfer one of the workers to another department, and equipment cost increased by $16 per hour while output increased by four carts per hour.

a) Compute the multifactor productivity (MFP) (labor plus equipment) under the Prior to buying the new equipment. The MFP (carts/$) = ________ (round to 4 decimal places).

b) Compute the % growth in productivity between the Prior to and after buying the new equipment. The growth in productivity = _________ (round to 2 decimal places)
Answer #1

Question: A Company that makes shopping carts for supermarkets and other stores recently purchased some new equipment that reduces the labor content of the jobs needed to produce the shopping carts. Prior to buying the new equipment, the company used 4 workers, who produced an average of 82 carts per hour. Workers receive $11 per hour, and machine coast was $52 per hour. With the new equipment, it was possible to transfer one of the workers to another department, and equipment cost increased by $16 per hour while output increased by four carts per hour.

Answer:

a. Compute the multifactor productivity (MFP) (labor plus equipment) under the Prior to buying the new equipment.

Multifactor Productivity (prior to buying new equipment):

Multifactor Productivity = 82 / (4 wkrs x 11 per hr) + 52 per hour

Multifactor Productivity = 82 / ((4 x 11) + 52)

Multifactor Productivity = 0.85416666666

Multifactor Productivity = 0.8542

b. Compute the % growth in productivity between the Prior to and after buying the new equipment.

We know that MFP of prior to buying the new equipment = 0.8542, now we need to find MFP after buying new equipment.

Multifactor Productivity (after buying new equipment):

Multifactor Productivity = 86 / (3 wkrs x 10 per hr) + 68 per hour

Multifactor Productivity = 86 / ((3 x 10) + 68)

Multifactor Productivity = 0.8775510204 = 0.8776

Now:

The growth in productivity:

Multifactor Productivity = ((0.8776 - 0.8542) / 0.8542) x 100

Multifactor Productivity = 2.7394052915

Multifactor Productivity = 2.74%